# RoboMP: A Robotic Multimodal Perception-Planning Framework with Multimodal Large Language Models

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) have shown impressive reasoning abilities and general intelligence in various domains. It inspires researchers to train end-to-end MLLMs or utilize large models to generate policies with human-selected prompts for embodied agents. However, these methods exhibit limited generalization capabilities on unseen tasks or scenarios, and overlook the multimodal environment information which is critical for robots to make decisions. In this paper, we introduce a novel Robotic Multimodal Perception-Planning (RoboMP) framework for robotic manipulation which consists of a Goal-Conditioned Multimodal Preceptor (GCMP) and a Retrieval-Augmented Multimodal Planner (RAMP). Specially, GCMP captures environment states by employing a tailored MLLMs for embodied agents with the abilities of semantic reasoning and localization. RAMP utilizes coarse-to-fine retrieval method to find the most-relevant policies as in-context demonstrations to enhance the planner. Extensive experiments demonstrate the superiority of RoboMP on both VIMA benchmark and real-world tasks, with around 10% improvement over the baselines.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) (GPT, 2023; Liu et al., 2023) have exhibited remarkable intelligence in various vision and language tasks. Inspired by the impressive performance, many efforts (Ahn et al., 2022; Huang et al., 2022) have been made on using large models to empower embodied agents with human-like intelligence. These approaches mainly use large models to train end-to-end policies (Driess et al., 2023; Brohan et al., 2023b, a) or generate plans with human-selected prompts (Liang et al., 2023; Huang et al., 2023b). However, end-to-end policies and human-selected prompts lack generalization and flexibility on unseen tasks.
Environment perception and task planning are two fundamental components for embodied agents to complete a task (Ahn et al., 2022; Sarch et al., 2023). Most of the existing work typically employs mainstream vision models as environment perceptors, such as YOLOv5 (Yue et al., 2022) and CLIP (Radford et al., 2021)). These models work well in simple scenarios where the categories of the objects are pre-defined or the relationships among objects are easy to be captured. However, they lack the capability to identify and locate objects in unseen scenarios or objects with intricate spatial relationships. For instance, as illustrated in Figure 1(a), the task entails picking up the “yellow block in the middle of the blue blocks”. The existing perceptors struggle to accurately identify and locate the specified yellow block since these models cannot understand the semantic information of the complex referring expression. Therefore, it necessitates the development of robot perceptors with multimodal understanding and reasoning capabilities. In light of this, we turn to MLLMs as environment perceptors since they have shown remarkable semantic understanding and vision perception abilities in various tasks (Liu et al., 2023). We accordingly propose a MLLM-based Goal-Conditioned Multimodal Perceptor (GCMP) with an advanced ability to detect objects with a given semantic-complex reference goal. As shown in Figure 1(a), GCMP enables robot agents to accurately identify and locate the referred object while the perceptors adopted in the existing embodied frameworks fail.
@@
In addition to the multimodal environment perceptor, the planning for subsequent execution is also critical. The existing policies mainly include end-to-end models (Brohan et al., 2023b) and prompt-based approaches (Huang et al., 2023b). The end-to-end policy integrates the perception and planning into a single model, thus requiring closed-loop robot data. However, in the real world, the closed-loop data is very limited due to the expensive human labor for collection. Consequently, these models are proven to overfitting (Ahn et al., 2022; Jiang et al., 2023) the data-collection scenario and show limited generalization in unseen environments or on new tasks. On the other hand, current prompt-based methods rely on manually designed and selected prompt templates to prompt LLMs to generate plans for a given task. They inherently lack generalization (Arenas et al., 2023; Huang et al., 2023a) in diversities of tasks that are highly different from the tasks given in the demonstrations of the prompt templates. To cover different types of tasks, these methods adopt large amounts of templates, which results in in-context attention distraction. Liu et al. (2021) have also demonstrated that redundant in-context examples fail to offer effective guidance, which is detrimental to the ability of the model (Liu et al., 2021). More importantly, these approaches (Huang et al., 2022; Vemprala et al., 2023) typically generate plans based solely on the text instruction, overlooking critical environmental information. For example, as shown in Figure 1(b), even for the same instruction, different environments lead to different plans. To address these issues, we propose a retrieval-augmented method which prompts MLLMs to generate plans via adaptively choosing the most relevant policies as demonstrations.
@@
In this paper, we aim to explore how to sufficiently leverage both multimodal information in an environment and the general intelligence in large models to enhance the perception and reasoning capabilities of embodied robots. We propose a novel Robotic Multimodal Perception-Planning framework (RoboMP) which is composed of a Goal-Conditioned Multimodal Perceptor (GCMP) and a Retrieval-Augmented Multimodal Planner (RAMP). Specifically, to endow the embodied model with semantic reasoning and localization capabilities, we propose GCMP to comprehensively perceive environmental information through the integration of a tailored MLLM. Meanwhile, to enhance the generalization of policy planning, we devise RAMP to adaptively find the most-relevant policies as in-context demonstrations with dedicated coarse retriever, fine reranker. Our main contributions are as followed:
@@
Extensive experiments show that our proposed RoboMP, which is composed of GCMP and RAMP, outperforms the baselines by around 10% on both VIMABench and real-world tasks.
@@ 2 Related Work
MLLMs for Robotic Manipulation. Robotic manipulation aims to complete a specific task by interacting with the environment. In recent, imitation learning (Brohan et al., 2023b, a; Chi et al., 2023) has achieved a great success in robotic learning. However, due to the task complexity, it needs large amounts of data to train a robot agent to achieve strong generalization capability (Fang et al., 2023a; Vuong et al., 2023; Shridhar et al., 2022), which is time-consuming and computational-unfriendly. Therefore, many efforts (Huang et al., 2023a, b; Xu et al., 2023) have been made on prompting LLMs to generate policies in a zero-shot manner to control robot agent. Liang et al. (2023) proposed using programs as the policy to control robots by generating calls to perception APIs and control APIs. It has been proven that the prompt is crucial when using LLMs to generate text in a zero-shot manner (Liu et al., 2021). Thus, many approaches (Arenas et al., 2023; Vemprala et al., 2023; Singh et al., 2023) explored the construction of the text prompt, including descriptions of the task details, robotic API libraries, environment feedback (Sarch et al., 2023). One constraint of previous studies is that they merely utilized the textual information to prompt LLMs, ignoring the significant multimodal information of the environment. Moreover, it would confuse the model if there are too many prompt templates in-context examples, resulting in a wrong execution plan. Thus, in order to alleviate these problems, we introduce a retrieval-augmented multimodal planner which sufficiently leverages the multimodal information and most relevant demonstrations for task planning.
@@
Retrieval Augmented Generation with MLLMs. Retrieval augmented generation (RAG) was first introduced to serve as more informative inputs to unlesh the extensive knowledge of LLMs (Lewis et al., 2021; Liu et al., 2021). Due to its effectiveness, it was subsequently introduced to the multimodal domain. It assists models generate answers by retrieving contents related to the original input as supplement context. RETRO (Borgeaud et al., 2022) introduced additional retrieval encoders to train a GPT model from scratch (Radford et al., 2019). Atlas (Izacard et al., 2023) incorporated similar retrieval encoders to continually finetune T5 model (Raffel et al., 2020). In addition, KNN-based retrieval methods (Fan et al., 2021) and cross modal semantic-based retrieval methods (Zhou et al., 2022) are commonly used to gather information from different modalities, aiming to generate more satisfied text by providing evidence. Considering the plans for robotic manipulation tasks are predominantly involved in the executed actions and target objects, we incorporate a task rewriting module. This module is introduced to extract essential textual information from task instructions. By integrating this module with a coarse-to-fine retrieval method, we adaptively select policies that exhibit semantic similarity as demonstrations, aiming to unveil the inherent capabilities of MLLMs.
@@ 3 Framework
In this section, we present a robotic multimodal perception-planning framework (RoboMP) that utilizes a Goal-Conditioned Multimodal Perceptor (GCMP) and a Retrieval-Augmented Multimodal Planner (RAMP) to enhance the perception and planning abilities of embdied agents. The framework of RoboMP is shown in Figure 2.
@@ 3.1.2 Training of Multimodal Perceptor
To learn a tailored multimodal perceptor for embodied agents, we formulate robotic data into instruction-tuning format, namely {image, ref_exp, coordinates} triplets, where the input consists of an image and a referential expression (ref_exp), and the output is the corresponding coordinates for manipulation. An example is shown in Figure 4. For model architecture, we adopt the ViT (Dosovitskiy et al., 2020) and flan-t5-xl (Chung et al., 2022) as our embodied vision encoder and language encoder. Then, we connect these two encoders through a MLP layer. In addition, we introduce a LoRA module (Hu et al., 2021) to tune the LLM efficiently and effectively.
@@ 3.2.2 Fine Reranker
In the plan generation stage, the order of the examples is crucial. Due to the design of positional encoding in the transformer architecture, researchers (Li et al., 2023; Rubin et al., 2021) have found that model generation tends to primarily focus on the content at the beginning and the end of the paragraph. Therefore, we introduce a reorder module with the aim of sorting the retrieved candidate code snippets. A straightforward solution is to simply concatenate it at the beginning of the whole candidate.
@@ 4.3 Implementation Details
For GCMP in our proposed RoboMP, we adopt the EVA-CLIP/g (Fang et al., 2023b) as the visual encoder of our perceptor. Regarding the LLM, we investigate the encoder-decoder architecture model flan-t5-xl (Chung et al., 2022). During training, the vision encoder and the LLM are frozen, only learning the weight of the fusion module and the LoRA module. We set the epoch to 10, the batch size to 128, the learning rates of the fusion module and LoRA module to 3e-5 and 1e-4, respectively. We adopt the AdamW optimizer and the cosine decay learning schedule. The overall training time is around 24 hours on a 8*A100-80G-SXM4 platform.
For RAMP in the proposed RoboMP, we adopt the GPT4/GPT3.5 and GPT4V as the text-only rewritter and the multimodal generator. The number of the most-relevant examples during coarse retriever and fine reranker is set to 5 and 2, respectively.
@@
Rotate | Rearrange |
Rearrange then restore |
@@
| RoboMP | 92.0 | 52.0 | 84.0 | 88.0 | 60.0 | 28.0 | 67.3 |
4.4 Results
Experimental Results on VIMABench.
The task difficulty of VIMABench ranges from L1 level to L4 level, with tasks at L1-L3 levels being seen, while tasks at L4 level are unseen. As shown in Table 1, our RoboMP outperforms other methods by a large extent, surpassing the VIMA baseline around 10% on the average performance of success ratio. This demonstrates the effectiveness of RoboMP for robotic manipulation.
Compared with the other baselines, the average performance improvements of RoboMP range from 4% to 8% on seen tasks (L1-L3), while a comprehensive improvement of approximately 20% on unseen tasks (L4). The reason is that end-to-end models tend to overfit to the training tasks due to the small amount of closed-loop data, resulting in limited generalization abilities on unseen tasks. The other prompted-based methods use fixed human-selected in-context examples in the prompt, thus lacking adaptability on diverse unseen tasks. In contrast, the proposed RoboMP shows a much better generalization ability on unseen tasks since RAMP in RoboMP can adaptively find the most relevant policies as their demonstrations to prompt the generator according to the task.
@@ 4.4 Results
We conduct experiments on 6 real-world tasks, comparing the performance of RoboMP with that of I2A. This is due to that both of them can be transferred from simulation to reality, without requiring to collect data for additional training. As illustrated in Table 2, our RoboMP outperforms I2A by 40% in terms of the overall average success ratio. It can be seen that I2A performs well in the task visual manipulation. However, it struggles to complete the manipulation tasks of objects with specific attributes such as same shape, let alone challenging tasks. In comparison, RoboMP shows consistent improvements across these two tasks.
It is noted that, on the tasks requiring to recognize objects with the same attribute or a specific relationship such as manipulate old neighbor, the performance of I2A drops significantly, even to zero. The reason is that the SAM+CLIP perceptor in I2A is unable to simultaneously identify multiple instances and is hardly handle the detection of objects with complex relationships. Benefiting from our proposed Goal-Conditioned Multimodal Perceptor (GCMP), RoboMP demonstrates a strong capability in dealing with such challenges.

### v2

# RoboMP2: A Robotic Multimodal Perception-Planning Framework with Multimodal Large Language Models

@@ Abstract
Multimodal Large Language Models (MLLMs) have shown impressive reasoning abilities and general intelligence in various domains. It inspires researchers to train end-to-end MLLMs or utilize large models to generate policies with human-selected prompts for embodied agents. However, these methods exhibit limited generalization capabilities on unseen tasks or scenarios, and overlook the multimodal environment information which is critical for robots to make decisions. In this paper, we introduce a novel Robotic Multimodal Perception-Planning (RoboMP2) framework for robotic manipulation which consists of a Goal-Conditioned Multimodal Preceptor (GCMP) and a Retrieval-Augmented Multimodal Planner (RAMP). Specially, GCMP captures environment states by employing a tailored MLLMs for embodied agents with the abilities of semantic reasoning and localization. RAMP utilizes coarse-to-fine retrieval method to find the most-relevant policies as in-context demonstrations to enhance the planner. Extensive experiments demonstrate the superiority of RoboMP2 on both VIMA benchmark and real-world tasks, with around 10% improvement over the baselines.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) (GPT, 2023; Liu et al., 2024) have exhibited remarkable intelligence in various vision and language tasks. Inspired by the impressive performance, many efforts (Ahn et al., 2022; Huang et al., 2022b) have been made on using large models to empower embodied agents with human-like intelligence. These approaches mainly use large models to train end-to-end policies (Driess et al., 2023; Brohan et al., 2023b, a) or generate plans with human-selected prompts (Liang et al., 2023; Huang et al., 2023b). However, end-to-end policies and human-selected prompts lack generalization and flexibility on unseen tasks.
Environment perception and task planning are two fundamental components for embodied agents to complete a task (Ahn et al., 2022; Sarch et al., 2023).
Most of the existing work typically employs mainstream vision models as environment perceptors, such as YOLOv5 (Yue et al., 2022) and CLIP (Radford et al., 2021)).
These models work well in simple scenarios where the categories of the objects are pre-defined or the relationships among objects are easy to be captured.
However, they lack the capability to identify and locate objects in unseen scenarios or objects with intricate spatial relationships.
For instance, as illustrated in Figure 1(a), the task entails picking up the “yellow block in the middle of the blue blocks”.
The existing perceptors struggle to accurately identify and locate the specified yellow block since these models cannot understand the semantic information of the complex referring expression.
Therefore, it necessitates the development of robot perceptors with multimodal understanding and reasoning capabilities.
In light of this, we turn to MLLMs as environment perceptors since they have shown remarkable semantic understanding and vision perception abilities in various tasks (Liu et al., 2024).
We accordingly propose a MLLM-based Goal-Conditioned Multimodal Perceptor (GCMP) with an advanced ability to detect objects with a given semantic-complex reference goal.
As shown in Figure 1(a), GCMP enables robots to accurately identify and locate referred objects while the per-
ceptors adopted in the existing embodied frameworks fail.
@@
In addition to the multimodal environment perceptor, the planning for subsequent execution is also critical. The existing policies mainly include end-to-end models (Brohan et al., 2023b) and prompt-based approaches (Huang et al., 2023b). The end-to-end policy integrates the perception and planning into a single model, thus requiring closed-loop robot data. However, in the real world, the closed-loop data is very limited due to the expensive human labor for collection. Consequently, these models are proven to overfitting (Ahn et al., 2022; Jiang et al., 2023) the data-collection scenario and show limited generalization in unseen environments or on new tasks. On the other hand, current prompt-based methods rely on manually designed and selected prompt templates to prompt LLMs to generate plans for a given task. They inherently lack generalization (Arenas et al., 2023; Huang et al., 2023a) in diversities of tasks that are highly different from the tasks given in the demonstrations of the prompt templates. To cover different types of tasks, these methods adopt large amounts of templates, which results in in-context attention distraction. Liu et al. (2022) have also demonstrated that redundant in-context examples fail to offer effective guidance, which is detrimental to the ability of the model (Liu et al., 2022). More importantly, these approaches (Huang et al., 2022b; Vemprala et al., 2023) typically generate plans based solely on the text instruction, overlooking critical environmental information. For example, as shown in Figure 1(b), even for the same instruction, different environments lead to different plans. To address these issues, we propose a retrieval-augmented method which prompts MLLMs to generate plans via adaptively choosing the most relevant policies as demonstrations.
@@
In this paper, we aim to explore how to sufficiently leverage both multimodal information in an environment and the general intelligence in large models to enhance the perception and reasoning capabilities of embodied robots. We propose a novel Robotic Multimodal Perception-Planning framework (RoboMP2) which is composed of a Goal-Conditioned Multimodal Perceptor (GCMP) and a Retrieval-Augmented Multimodal Planner (RAMP). Specifically, to endow the embodied model with semantic reasoning and localization capabilities, we propose GCMP to comprehensively perceive environmental information through the integration of a tailored MLLM. Meanwhile, to enhance the generalization of policy planning, we devise RAMP to adaptively find the most-relevant policies as in-context demonstrations with dedicated coarse retriever, fine reranker. Our main contributions are as followed:
@@
Extensive experiments show that our proposed RoboMP2, which is composed of GCMP and RAMP, outperforms the baselines by around 10% on both VIMABench and real-world tasks.
@@ 2 Related Work
MLLMs for Robotic Manipulation. Robotic manipulation aims to complete a specific task by interacting with the environment. In recent, imitation learning (Brohan et al., 2023b, a; Chi et al., 2023) has achieved a great success in robotic learning. However, due to the task complexity, it needs large amounts of data to train a robot agent to achieve strong generalization capability (Fang et al., 2023a; Vuong et al., 2023; Shridhar et al., 2022), which is time-consuming and computational-unfriendly. Therefore, many efforts (Huang et al., 2023a, b; Xu et al., 2023) have been made on prompting LLMs to generate policies in a zero-shot manner to control robot agent. Liang et al. (2023) proposed using programs as the policy to control robots by generating calls to perception APIs and control APIs. It has been proven that the prompt is crucial when using LLMs to generate text in a zero-shot manner (Liu et al., 2022). Thus, many approaches (Arenas et al., 2023; Vemprala et al., 2023; Singh et al., 2023) explored the construction of the text prompt, including descriptions of the task details, robotic API libraries, environment feedback (Sarch et al., 2023). One constraint of previous studies is that they merely utilized the textual information to prompt LLMs, ignoring the significant multimodal information of the environment. Moreover, it would confuse the model if there are too many prompt templates in-context examples, resulting in a wrong execution plan. Thus, in order to alleviate these problems, we introduce a retrieval-augmented multimodal planner which sufficiently leverages the multimodal information and most relevant demonstrations for task planning.
@@
Retrieval Augmented Generation with MLLMs. Retrieval augmented generation (RAG) was first introduced to serve as more informative inputs to unlesh the extensive knowledge of LLMs (Lewis et al., 2020; Liu et al., 2022). Due to its effectiveness, it was subsequently introduced to the multimodal domain. It assists models generate answers by retrieving contents related to the original input as supplement context. RETRO (Borgeaud et al., 2022) introduced additional retrieval encoders to train a GPT model from scratch (Radford et al., 2019). Atlas (Izacard et al., 2023) incorporated similar retrieval encoders to continually finetune T5 model (Raffel et al., 2020). In addition, KNN-based retrieval methods (Fan et al., 2021) and cross modal semantic-based retrieval methods (Zhou et al., 2022) are commonly used to gather information from different modalities, aiming to generate more satisfied text by providing evidence. Considering the plans for robotic manipulation tasks are predominantly involved in the executed actions and target objects, we incorporate a task rewriting module. This module is introduced to extract essential textual information from task instructions. By integrating this module with a coarse-to-fine retrieval method, we adaptively select policies that exhibit semantic similarity as demonstrations, aiming to unveil the inherent capabilities of MLLMs.
Task and Motion Planning. It (Garrett et al., 2021; Silver et al., 2023; Kaelbling & Lozano-Pérez, 2011) selects the sequence of high-level actions (Fikes & Nilsson, 1971; Nau et al., 1999) that the robot should take, the hybrid parameter values that determine how the action is performed, and the low-level motions (LaValle, 2006) that safely execute the action. Traditionally, these approaches build on research through optimization (Toussaint, 2015) or symbolic reasoning (Kumar et al., 2022), but more recently machine learning has been applied to aspects of the problem via learned representations (Silver et al., 2023; Eysenbach et al., 2019; Xu et al., 2018, 2019; Shah et al., 2021), learned task primitives, and more. Some works utilize language for planning and grounding (Huang et al., 2022a, b; Ahn et al., 2022; Kollar et al., 2010; Tellex et al., 2011). Others have approached the problem through hierarchical learning (Nair & Finn, 2019; Xia et al., 2020; Shah et al., 2021; Hafner et al., 2022). In this work, we leverage MLLMs and their semantic knowledge to find feasible plans.
@@ 3 Framework
In this section, we present a robotic multimodal perception-planning framework (RoboMP2) that utilizes a Goal-Conditioned Multimodal Perceptor (GCMP) and a Retrieval-Augmented Multimodal Planner (RAMP) to enhance the perception and planning abilities of embdied agents. The framework of RoboMP2 is shown in Figure 2.
@@ 3.1.2 Training of Multimodal Perceptor
To learn a tailored multimodal perceptor for embodied agents, we formulate robotic data into instruction-tuning format, namely {image, ref_exp, coordinates} triplets, where the input consists of an image and a referential expression (ref_exp), and the output is the corresponding coordinates for manipulation. An example is shown in Figure 4. For model architecture, we adopt the ViT (Dosovitskiy et al., 2020) and flan-t5-xl (Chung et al., 2024) as our embodied vision encoder and language encoder. Then, we connect these two encoders through a MLP layer. In addition, we introduce a LoRA module (Hu et al., 2021) to tune the LLM efficiently and effectively.
@@ 3.2.2 Fine Reranker
In the plan generation stage, the order of the examples is crucial. Due to the design of positional encoding in the transformer architecture, researchers (Li et al., 2023; Rubin et al., 2022) have found that model generation tends to primarily focus on the content at the beginning and the end of the paragraph. Therefore, we introduce a reorder module with the aim of sorting the retrieved candidate code snippets. A straightforward solution is to simply concatenate it at the beginning of the whole candidate.
@@
| L1 | L2 | L3 | L4 | Avg. | |
| End-to-end models | |||||
| Gato† | 58.1 | 53.2 | 43.5 | 12.4 | 41.8 |
| Flamingo† | 47.5 | 46.0 | 40.8 | 12.1 | 36.6 |
| VIMA† | 81.6 | 81.5 | 79.0 | 48.9 | 72.7 |
| RT-2 | 72.8 | 70.3 | 66.8 | 47.0 | 64.2 |
| MLLM Planners | |||||
| CaP | 71.2 | 70.0 | 42.8 | 44.7 | 57.2 |
| VisualProg | 49.7 | 47.7 | 69.9 | 52.2 | 54.9 |
| I2A† | 77.0 | 76.2 | 66.6 | 49.8 | 65.0 |
| RoboMP2 | 89.0 (+7.4) | 85.9 (+4.4) | 86.8 (+7.8) | 68.0 (+19.1) | 82.4 (+9.7) |
| Task Name | I2A | RoboMP2 |
| Basic Task | ||
| Visual Manipulation | 85.0 | 90.0 |
| Same Shape | 45.0 | 90.0 |
| Same Color | 40.0 | 90.0 |
| Challenging Task | ||
| Manipulate Old Neighbor | 0.0 | 70.0 |
| Pick in Order then Restore | 65.0 | 80.0 |
| Interfering Manipulation | 0.0 | 55.0 |
| Avg. Success Ratio (%) | 39.2 | 79.2 |
@@ 4.3 Implementation Details
For GCMP in our proposed RoboMP2, we adopt the EVA-CLIP/g (Fang et al., 2023b) as the visual encoder of our perceptor. Regarding the LLM, we investigate the encoder-decoder architecture model flan-t5-xl (Chung et al., 2024). During training, the vision encoder and the LLM are frozen, only learning the weight of the fusion module and the LoRA module. We set the epoch to 10, the batch size to 128, the learning rates of the fusion module and LoRA module to 3e-5 and 1e-4, respectively. We adopt the AdamW optimizer and the cosine decay learning schedule. The overall training time is around 24 hours on a 8*A100-80G-SXM4 platform.
For RAMP in the proposed RoboMP2, we adopt the GPT4/GPT3.5 and GPT4V as the text-only rewritter and the multimodal generator. The number of the most-relevant examples during coarse retriever and fine reranker is set to 5 and 2, respectively.
4.4 Results
Experimental Results on VIMABench.
The task difficulty of VIMABench ranges from L1 level to L4 level, with tasks at L1-L3 levels being seen, while tasks at L4 level are unseen. As shown in Table 1, our RoboMP2 outperforms other methods by a large extent, surpassing the VIMA baseline around 10% on the average performance of success ratio. This demonstrates the effectiveness of RoboMP2 for robotic manipulation.
Compared with the other baselines, the average performance improvements of RoboMP2 range from 4% to 8% on seen tasks (L1-L3), while a comprehensive improvement of approximately 20% on unseen tasks (L4). The reason is that end-to-end models tend to overfit to the training tasks due to the small amount of closed-loop data, resulting in limited generalization abilities on unseen tasks. The other prompted-based methods use fixed human-selected in-context examples in the prompt, thus lacking adaptability on diverse unseen tasks. In contrast, the proposed RoboMP2 shows a much better generalization ability on unseen tasks since RAMP in RoboMP2 can adaptively find the most relevant policies as their demonstrations to prompt the generator according to the task.
@@ 4.4 Results
manipulate |
Rotate | Rearrange | Rearrange then restore |
@@
| RoboMP2 | 92.0 | 52.0 | 84.0 | 88.0 | 60.0 | 28.0 | 67.3 |
@@
We conduct experiments on 6 real-world tasks, comparing the performance of RoboMP2 with that of I2A. This is due to that both of them can be transferred from simulation to reality, without requiring to collect data for additional training. As illustrated in Table 2, our RoboMP2 outperforms I2A by 40% in terms of the overall average success ratio. It can be seen that I2A performs well in the task visual manipulation. However, it struggles to complete the manipulation tasks of objects with specific attributes such as same shape, let alone challenging tasks. In comparison, RoboMP2 shows consistent improvements across these two tasks.
It is noted that, on the tasks requiring to recognize objects with the same attribute or a specific relationship such as manipulate old neighbor, the performance of I2A drops significantly, even to zero. The reason is that the SAM+CLIP perceptor in I2A is unable to simultaneously identify multiple instances and is hardly handle the detection of objects with complex relationships. Benefiting from our proposed Goal-Conditioned Multimodal Perceptor (GCMP), RoboMP2 demonstrates a strong capability in dealing with such challenges.